### v1

# UniRGB-IR: A Unified Framework for Visible

Abstract
Semantic analysis on visible (RGB) and infrared (IR) images has gained attention for its ability to be more accurate and robust under low-illumination and complex weather conditions. Due to the lack of pre-trained foundation models on the large-scale infrared image datasets, existing methods prefer to design task-specific frameworks and directly fine-tune them with pre-trained foundation models on their RGB-IR semantic relevance datasets, which results in poor scalability and limited generalization. In this work, we propose a scalable and efficient framework called UniRGB-IR to unify RGB-IR downstream tasks, in which a novel adapter is developed to efficiently introduce richer RGB-IR features into the pre-trained RGB-based foundation model. Specifically, our framework consists of a vision transformer (ViT) foundation model, a Multi-modal Feature Pool (MFP) module and a Supplementary Feature Injector (SFI) module. The MFP and SFI modules cooperate with each other as an adpater to effectively complement the ViT features with the contextual multi-scale features. During training process, we freeze the entire foundation model to inherit prior knowledge and only optimize the MFP and SFI modules. Furthermore, to verify the effectiveness of our framework, we utilize the ViT-Base as the pre-trained foundation model to perform extensive experiments. Experimental results on various RGB-IR downstream tasks demonstrate that our method can achieve state-of-the-art performance. The source code and results are available at https://github.com/PoTsui99/UniRGB-IR.git
Index Terms:
RGB-IR Downstream Tasks, Multi-modal Fusion, Adapter, Transformer.
I Introduction
Single visible (RGB) image semantic analysis is a common practice in computer vision and has been widely used in a variety of vision tasks [1, 2, 3]. With the development of physical devices, more and more general-purpose foundation backbones [4, 5] pre-trained on a large-scale dataset (e.g., ImageNet [6] and COCO [7]) have been designed for various downstream tasks. Since these pre-trained foundation models implicitly encode a large amount of prior knowledge, many researchers [8, 9] always fine-tune them on their semantic relevance datasets to significantly improve the performance on downstream tasks and speed up training convergence.
Figure 1: Existing full fine-tuning methods vs. our UniRGB-IR framework. (a) Existing methods use pre-trained RGB-based foundation models and fully fine-tune them on their RGB-IR semantic relevance datasets. (b) We utilize the Adpater [10] to propose a new unified framework called UniRGB-IR, which can efficiently introduce richer RGB-IR features into the pre-trained foundation model for downstream tasks.
However, visible cameras have proved to be difficult in providing reliable imaging [11, 12], due to their limited bandwidth across the spectrum. Especially in low-illumination and complex weather conditions, the useful information of RGB images is relatively scarce. Therefore, infrared (IR) modality with stronger low-light adaptability are used as additional information to supplement RGB modality [11, 13]. Afterwards, the joint use of RGB and IR images has been applied in more and more semantic analysis tasks.
Due to the lack of pre-trained foundation models on the large-scale infrared image datasets, most existing methods [14, 15] prevailing trend to use pre-trained RGB baselines and fine-tune them on their RGB-IR semantic relevance datasets as shown in Figure 1 (a). For example, RGB-IR object detectors [14, 16] utilize RGB-based models as strong baselines and fine-tune them to extract RGB and IR features. SwinNet [17] is explored based on the RGB-based transformer to extract hierarchical features of each modality for the detection of salient RGB-IR objects. Similarly, Ha et al. [18] propose a new baseline by aggregating infrared modality features into the RGB-based framework and retrain it in the RGB-IR semantic segmentation benchmark. Among the efforts to achieve competitive results, these approaches still suffer from two major issues. (1) Different downstream tasks require customized model structures, resulting in poor model scalability and excessive parameter storage. (2) Full fine-tuning strategy impairs the prior knowledge encoded in the foundation model, which limits the generalization potential of the fine-tuned model. Based on these issues, a question that naturally arises is whether we can adapt the RGB-based foundation model to efficiently construct a unified framework for RGB-IR downstream tasks.
To achieve this purpose, in this paper, we explore an efficient and scalable framework for RGB-IR downstream tasks and, for the first time, propose a Unified model for RGB-IR downstream tasks called UniRGB-IR as shown in Figure 1 (b). Drawing inspiration from recent advances in adapters [10, 19], which is initially used to yield a extensible model to effectively exploit the potential representation of foundation models in natural language processing (NLP) field, we aim to develop an adapter to dynamically introduce richer RGB-IR features into the pre-trained foundation model through an additional CNN branch. Therefore, without modifying the original foundation model, the powerful pre-trained weights can be directly loaded and frozen to accelerate training convergence and achieve efficient fine-tuning on downstream tasks. Specifically, we design two core modules in our UniRGB-IR: a Multi-modal Feature Pool module (MFP) and a Supplementary Feature Injector module (SFI). In the MFP module, the multi-receptive field convolution and the feature pyramid are deployed to capture contextual multi-scale features from RGB and IR images. These features are fused at different scales to complement foundation models for different RGB-IR downstream tasks. For the SFI module, a cross-attention mechanism is used to dynamically inject the required features into the pre-trained foundation model, which facilitates our UniRGB-IR framework with a robust feature representation ability. To inherit prior knowledge of foundation model pre-trained on large-scale datasets, we utilize the adapter tuning paradigm instead of the full fine-tuning manner. In specific, we freeze the entire pre-trained weights and only optimize the MFP module and the SFI module. Therefore, our framework can be utilized as a unified framwork to efficiently fine-tune on various RGB-IR downstream tasks.
Overall, our contributions are summarized as follows:
  • We explore an efficient and scalable framework called UniRGB-IR to unify RGB-IR downstream tasks. To the best of our knowledge, this is the first attempt to construct a unified framework for RGB-IR downstream tasks.
  • We design a Multi-modal Feature Pool module and a Supplementary Feature Injector module. The former extracts contextual multi-scale features from two modality images, and the latter dynamically injects the required features into the pre-trained model. These two modules can be efficiently fine-tuned with adapter tuning paradigm to complement the pre-trained foundation model with richer RGB-IR features for RGB-IR downstream tasks.
  • We incorporate the vision transformer foundation model into the UniRGB-IR framework to evaluate the effectiveness of our method on RGB-IR downstream tasks, including RGB-IR object detection, RGB-IR semantic segmentation, and RGB-IR salient object detection. Extensive experimental results demonstrate that our methods can efficiently achieve superior performance on these RGB-IR downstream tasks.
II Related Work
II-A Vision Foundation Models
Recently, thanks to the powerful long-distance modeling capability, vision transformers (ViT) [4] have been widely used as the foundation model in many vision tasks to achieve competitive results. The original ViT is a plain, non-hierarchical structure for image classification. Based on it, ViTDet [20] also constructs a non-hierarchical model by incorporating the feature pyramid. However, the non-hierarchical structure lacks rich feature representation, resulting in unsatisfactory performance. Subsequently, various hierarchical transformers [21, 22, 23, 5] have been proposed for different downstream vision tasks. Swin Transformer [21] and Multiscale Vision Transformer [22] are designed based on the ViT model to explore multi-scale features to improve the performance of image classification and object detection tasks. Besides, PVT [24] performs global attention on the downsampled key and value maps for dense prediction. In this paper, we also introduce the ViT model as the pre-trained foundation model to construct a unified framework for different RGB-IR downstream tasks.
II-B RGB-IR Downstream Tasks
Since infrared cameras have stronger low-light adaptability to capture object thermal radiation, infrared images have been widely used together with RGB images in various RGB-IR downstream tasks in recent years.
RGB-IR Object Detection. Zhang et al. [25] explore a two-stream SSD [26] structure to capture the contextual enhanced features for RGB-IR object detection. Besides, AR-CNN [27] is presented based on faster R-CNN [28] to align RGB and IR features. With the emergence of transformers, Yuan et al. [29] propose a complementary fusion transformer (CFT) module to achieve advanced detection results. Furthermore, C2Former [12] is a novel transformer block and can be incorporated into exited pre-trained models to increase intra- and inter-modality feature representations.
RGB-IR Semantic Segmentation. MFNet [18] is proposed to incorporate infrared features into the RGB-based framework to perform RGB-IR semantic segmentation. Based on the transformer structure, Wu et al. [30] propose a novel CCFFNet to excavate discriminative and complementary modality features for RGB-IR semantic segmentation. Moreover, CMX [31] is designed as a universal cross-modal fusion framework for RGB-IR semantic segmentation in an interactive fusion manner.
Figure 2: The overall architecture of our UniRGB-IR. In our framework, a ViT model with different numbers of ViT blocks is deployed as a foundation model, which is divided into $N$ (usually $N=4$) stages for feature interaction. Besides, a Multi-modal Feature Pool (MFP) is designed to extract contextual multi-scale RGB-IR features and a Supplementary Feature Injector (SFI) is explored to introduce these features into the ViT foundation model. During training, we freeze the entire ViT model weights and only optimize the MFP and SFI modules.
RGB-IR Salient Object Detection. SwinNet [17] is designed based on the Swin Transformer to extract hierarchical information of each modality, which achieves impressive results. CAVER [32] introduces the transformer to rethink the bi-modal salient object detection from a sequence-to-sequence perspective, which increases the model interpretability. Recently, Zhou et al. [33] transfer a large amount of knowledge learned in the transformer-based network to lightweight WaveNet through the distillation method.
The above methods attempt to design task-oriented structures to improve performance on corresponding downstream tasks. They either train the designed model from scratch or adopt a full fine-tuning strategy on a pre-trained model. Unlike these methods, we take a further step and deploy an adapter based on the pre-trained foundation model to unify the RGB-IR downstream tasks.
II-C Adapter
In the NLP field, Adapter [10] first fixes the original foundation backbone and introduces new modules into the transformer for task-specific fine-tuning, thereby effectively adapting the pre-trained backbone to downstream NLP tasks. Afterwards, Adapter has been widely studied in computer vision. ViT-Adapter [34] and Low-Rank Adapter [35] are designed to introduce a modest number of trainable parameters into the ViT and fine-tuning efficiently for dense prediction tasks. In addition, PC-Adapter [36] explores an attention-based adapter to preserve global shape knowledge for domain adaptation on point cloud data. Recently, Adapter is also used as a parameter-efficient training technology for vision-and-language tasks [37, 38].
In this paper, we aim to explore an adapter capable of converting the IR features into the RGB features to fit the pre-trained foundation model, which remains challenging to design this unified framework. Our UniRGB-IR is the first to propose utilizing adapter to unify RGB-IR downstream tasks.
III Method
III-A Overall Architecture
The overall framework of UniRGB-IR is illustrated in Figure 2, which consists of three parts: vision transformer model, Multi-modal Feature Pool (MFP) module, and Supplementary Feature Injector (SFI) module. In our framework, the ViT model is utilized as the pre-trained foundation model and frozen during the training process. Specifically, for the ViT model, the RGB image is directly fed into the patch embedding process to obtain the $D$-dimensional feature tokens, which are usually 1/16 of the original image resolution. To complement the richer features required for various RGB-IR downstream tasks, we feed the RGB and IR images into the MFP module to extract contextual multi-scale features from two modalities (eg. 1/8, 1/16 and 1/32 of the original image resolution). Afterwards, these richer features are dynamically injected into the features of ViT model through the SFI module, which can adaptively introduce the required RGB-IR features into the ViT model. To fully integrate the extracted features into the ViT model, we add the SFI module at the beginning of each stage. Thus, after $N$ stages of feature injection, the final features from ViT model can be leveraged for various RGB-IR downstream tasks.
Figure 3: Structure of the multi-modal feature pool (MFP) module. We explore the multiple perceptions to expand the receptive field of contextual feature extraction and utilize the feature pyramid to obtain the multi-scale features. In this module, Squeeze-Excitation(SE) attention is utilized to fuse two modality features. Thus, the contextual multi-scale features can be provided for the various RGB-IR downstream tasks.
III-B Multi-modal Feature Pool
To complement the rich feature representations for different RGB-IR downstream tasks, we introduce a multi-modal feature pool (MFP) module, including multiple perception and feature pyramid. The former can extract the contextual features with different convolution kernels, which achieve the long-distance modeling capability of CNNs. Different from existing works [39, 40] that increase the width or depth of the model, we efficiently achieve multi-receptive field perception in the channel dimension. As for feature pyramid, it can obtain multi-scale features to enhance the small object features. Thus, these two operations are connected in series to construct MFP module, as shown in Figure 3.
Specifically, for the input RGB ($H\times W\times 3$) and IR ($H\times W$) images, we first employ a stem block borrowed from ResNet [41] to extract two modality features $F^{rgb}_{1}$ and $F^{ir}_{1}\in\mathbb{R}^{H/4\times W/4\times C}$. Then, these two features are split into four equal parts by utilizing the channel splitting technique [42]. To extract multi-receptive field perception, each part is subjected to convolution operations with different kernel sizes ($3\times 3$, $3\times 3$, $5\times 5$ and $7\times 7$). Then, we fuse each processed feature from two modalities by using a SE attention module (shown in Figure 3). Therefore, we concatenate each fused part to obtain RGB-IR contextual features $F_{fus}$. The above process can be represented as:
$F_{fus}=\varGamma_{k=1}^{4}(Fusion(W_{k}^{rgb}*f_{k}^{rgb},W_{k}^{ir}*f_{k}^{ir})),$ (1)
where $F_{fus}\in\mathbb{R}^{H/4\times W/4\times C}$, $f_{k}^{rgb}$ and $f_{k}^{ir}$ are the $k$-th part of $F_{1}^{rgb}$ and $F_{1}^{ir}$ features respectively, $W_{k}$ is the convolution with $k$-th kernel size, $\varGamma$ is the concatenation operation.
For the feature pyramid, a stack of three $3\times 3$ convolutions with $stride=2$ is conducted to downsample the size of the feature maps. Then, features of each scale are fed into a $1\times 1$ convolution to project the feature maps to $D$ dimensions. Therefore, we can obtain a set of multi-scale features {$F_{2}$, $F_{3}$, $F_{4}$} with 1/8, 1/16, and 1/32 of the original image resolution, respectively. Finally, we flatten and concatenate these features into feature tokens $F_{mfp}\in\mathbb{R}^{(\frac{HW}{8^{2}}+\frac{HW}{16^{2}}+\frac{HW}{32^{2}})\times D}$, which will be used as supplementary features for the ViT model.
Figure 4: Structure of supplementary feature injector (SFI) module. A gated fusion unit (GFU) is utilized to dynamically fuse the current and the last injected features.
III-C Supplementary Feature Injector
As shown in Figure 4, we propose a supplementary feature injector (SFI) module to adaptively introduce the contextual multi-scale features without altering the ViT structure. Since the sequence lengths of contextual multi-scale features $F_{mfp}$ and the ViT features $F_{vit}^{i}$ are different, to address this, we employ sparse attention (eg. [43] and [44]) to dynamically sample supplementary features from each scale. To be specific, we utilize the ViT features $F_{vit}^{i}\in\mathbb{R}^{\frac{HW}{16^{2}}\times D}$ as the query, and the contextual multi-scale features $F_{mfp}\in\mathbb{R}^{(\frac{HW}{8^{2}}+\frac{HW}{16^{2}}+\frac{HW}{32^{2}})\times D}$ as the key and value:
$\tilde{F}_{sfi}^{i}=Attention(LN(F_{vit}^{i}),LN(F_{mfp})),$ (2)
where $Attention(\cdot)$ is the attention layer and $LN(\cdot)$ is LayerNorm [45] which is intended to reduce modality discrenpancy. Furthermore, we adopt progressive injection to introduce contextual multi-scale features, which can balance the foundation model features and the injected features $F_{sfi}^{i}$. Thus, a gated fusion unit is explored to predict the fusion weight $z$ to gate $F_{sfi}^{i-1}$ and $\tilde{F}_{sfi}^{i}$ for dynamic fusion. Specifically, we concatenate the two features $F_{sfi}^{i-1}$ and $\tilde{F}_{sfi}^{i}$ and feed it into linear layer to predict the weight $z$. Then, $z$ and $1-z$ are used to fuse $F_{sfi}^{i-1}$ and $\tilde{F}_{sfi}^{i}$ features respectively. The final output features $F_{sfi}^{i}$ of SFI module can be formulated as:
$F_{sfi}^{i}=\begin{cases}\tilde{F}_{sfi}^{i},&i=1\\ {GFU}(\tilde{F}_{sfi}^{i},F_{sfi}^{i-1}),&i=2\ldots N\end{cases}$ (3)
where $GFU(\cdot,\cdot)$ is the gated fusion unit shown in Figure 4.
TABLE I: Comparison results (mAP, in%) on the FLIR and LLVIP dataset. The best results are highlighted in red and the second-best are highlighted in blue. “–” indicates that the author did not provide the corresponding results.
FLIR LLVIP
Methods Modality mAP $\text{mAP}_{50}$ $\text{mAP}_{75}$ mAP $\text{mAP}_{50}$ $\text{mAP}_{75}$
SSD [26] IR 24.6 57.0 18.0 53.5 90.2 57.9
RetinaNet [46] IR 31.5 66.1 25.3 55.1 94.8 57.6
Faster R-CNN [28] IR 37.6 75.8 31.6 54.5 94.6 57.6
Cascade R-CNN (ViT-B) [47] IR 41.9 78.6 37.3 60.4 94.9 67.5
DDQ-DETR [48] IR 37.1 73.9 32.2 58.6 93.9 64.6
SSD [26] RGB 18.8 46.3 13.1 39.8 82.6 31.8
RetinaNet [46] RGB 21.9 51.2 15.2 42.8 88.0 34.4
Faster R-CNN [28] RGB 27.7 62.2 21.2 45.1 87.0 41.2
Cascade R-CNN (ViT-B) [47] RGB 33.3 69.3 26.2 51.7 90.3 54.7
DDQ-DETR [48] RGB 30.9 64.9 24.5 46.7 86.1 45.8
GAFF [49] RGB+IR 37.4 74.6 31.3 55.8 94.0 60.2
ProbEn [50] RGB+IR 37.9 75.5 31.8 51.5 93.4 50.2
LGADet [51] RGB+IR - 74.5 - - - -
CSAA [52] RGB+IR 41.3 79.2 37.4 59.2 94.3 66.6
UniRGB-IR (Ours) RGB+IR 44.1 81.4 40.2 63.2 96.1 72.2
TABLE II: Comparison results ($\text{MR}^{\text{-2}}$, in%) under ‘All-dataset’ settings of different pedestrian distances, occlusion levels, and light conditions (Day and Night) on the KAIST dataset. The best and second results are highlighted in red and blue respectively.
Methods Near Medium Far None Partial Heavy Day Night All
ACF [53] 28.74 53.67 88.20 62.94 81.40 88.08 64.31 75.06 67.74
Halfway Fusion [54] 8.13 30.34 75.70 43.13 65.21 74.36 47.58 52.35 49.18
FusionRPN+BF [55] 0.04 30.87 88.86 47.45 56.10 72.20 52.33 51.09 51.70
IAF R-CNN [56] 0.96 25.54 77.84 40.17 48.40 69.76 42.46 47.70 44.23
IATDNN+IASS [57] 0.04 28.55 83.42 45.43 46.25 64.57 49.02 49.37 48.96
CIAN [25] 3.71 19.04 55.82 30.31 41.57 62.48 36.02 32.38 35.53
MSDS-R-CNN [58] 1.29 16.19 63.73 29.86 38.71 63.37 32.06 38.83 34.15
AR-CNN [27] 0.00 16.08 69.00 31.40 38.63 55.73 34.36 36.12 34.95
MBNet [59] 0.00 16.07 55.99 27.74 35.43 59.14 32.37 30.95 31.87
TSFADet [16] 0.00 15.99 50.71 25.63 37.29 65.67 31.76 27.44 30.74
CMPD [60] 0.00 12.99 51.22 24.04 33.88 59.37 28.30 30.56 28.98
CAGTDet [29] 0.00 14.00 49.40 24.48 33.20 59.35 28.79 27.73 28.96
C2Former [12] 0.00 13.71 48.14 23.91 32.84 57.81 28.48 26.67 28.39
UniRGB-IR (Ours) 0.00 13.44 38.21 20.26 31.67 55.03 25.93 23.95 25.21
III-D Adapter Tuning Paradigm
To fully inherit the prior knowledge of the ViT pre-trained on large-scale datasets, we explore the adapter tuning paradigm instead of the full fine-tuning manner. For the dataset $D=\{(x_{j},{gt}_{j})\}_{j=1}^{M}$ of the downstream task, full fine-tuning process calculates the loss between the prediction and the ground truth, which can be formulated as:
$\mathcal{L}(D,\theta)=\sum_{j=1}^{M}\operatorname{loss}(F_{\theta}(x_{j}),{gt}_{j}),$ (4)
where loss represents the loss function and $F_{\theta}$ denotes the entire network parameterized by $\theta$. Afterwards, $\theta$ is optimized through the formula:
[…]
However, in our adapter tuning paradigm, the parameter $\theta$ consists of two parts, one part is the parameter in the original ViT model $\theta_{V}$, and the other part is the parameter in our UniRGB-IR $\theta_{A}$. During training, we freeze the parameter $\theta_{V}$ and only optimize the parameter $\theta_{A}$. Thus, the loss function and optimization of our a can be represented as:
$\mathcal{L}(D,{\theta_{V}},{\theta_{A}})=\sum_{j=1}^{M}\operatorname{loss}(F_{\theta_{V},\theta_{A}}(x_{j}),{gt}_{j}),$ (6)
$\theta_{A}\leftarrow\underset{\theta_{A}}{\arg}\min\mathcal{L}(D,\theta_{V},\theta_{A}).$ (7)
TABLE III: Results on the PST900 dataset. The best results are highlighted in red and the second-best are highlighted in blue.
“–” indicates that the author did not provide the corresponding results.
Methods Background Fire-Extinguisher Backpack Hand-Drill Survivor mAcc mIoU
Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU
MFNet [18] 99.9 98.7 71.8 67.4 52.8 52.4 46.7 39.3 18.8 18.9 58.0 55.3
RTFNet [61] 99.9 99.0 78.6 54.8 62.5 60.8 76.7 61.0 65.2 62.5 76.6 67.6
CCNet [62] 99.6 98.7 88.1 73.8 76.0 73.0 54.1 51.0 49.5 33.5 73.46 66.0
PSTNet [63] - 98.9 - 70.1 - 69.2 - 53.6 - 50.0 - 68.4
ACNet [64] 99.8 99.3 84.9 60.0 85.6 83.2 53.6 51.5 69.1 65.2 78.7 71.8
FDCNet [65] 99.7 99.2 91.8 71.5 77.5 72.2 82.5 70.4 78.4 72.4 86.0 77.1
CGFNet [66] 99.7 99.3 96.3 71.7 87.6 82.0 70.5 59.7 86.3 77.4 88.1 78.0
EGFNet [67] 99.5 99.2 95.2 71.3 94.2 83.1 98.0 64.7 83.3 74.3 94.0 78.5
CCFFNet [30] 99.9 99.4 87.0 79.3 80.2 77.8 88.3 77.4 77.9 73.4 86.7 81.5
RSFNet [68] - 99.4 - 75.4 - 84.9 - 72.9 - 70.1 - 80.5
SGFNet [69] 99.8 99.4 89.4 75.6 90.4 85.4 94.0 76.7 82.7 76.7 91.2 82.8
UniRGB-IR (Ours) 99.7 99.5 97.0 72.0 93.3 87.7 95.5 78.0 86.1 77.8 94.3 82.8
IV Experiments
To evaluate the effectiveness of our UniRGB-IR, we utilize the ViT-Base model (pre-trained on COCO [7] dataset) as the foundation model and utilize this framework to perform different RGB-IR downstream tasks. During training, we freeze the ViT-Base model and only optimize the MFP and SFI modules. We evaluate and compare our method with various competitive models, including CNN-based and Transformer-based models. Besides, our evaluation spans various tasks, including RGB-IR object detection on FLIR [70], LLVIP [71], and KAIST [53] datasets, RGB-IR semantic segmentation on MFNet [18] and PST9000 [63] datasets and RGB-IR salient object detection (see supplementary materials). Furthermore, ablation experiments on the designed modules and qualitative experiments are also conducted to verify that the UniRGB-IR framework can be leveraged as a unified framework to efficiently introduce IR features into the foundation model to achieve superior performance.
Figure 5: The Miss rate curves for all detectors are plotted against False Positives Per Image (FPPI) under the IoU threshold of 0.5 on the KAIST dataset.
IV-A RGB-IR Object Detection
Datasets. Our object detection experiments are based on the three paired RGB and IR object detection datasets. FLIR [70] is a paired visible and infrared object detection dataset, including daytime and night scenes, which has 4,129 aligned RGB-IR image pairs for training and 1,013 for testing. For LLVIP [71] dataset, it contains 15,488 aligned RGB-IR image pairs, of which 12,025 images are used for training and 3,463 images for testing. As for KAIST [53] dataset, it is a aligned multispectral pedestrian deteciton dataset, in which 8,963 and 2,252 image pairs are utilized for training and testing.
Metrics. For FLIR and LLVIP datasets, we employ mean Average Precision (mAP) to evaluate the detection performance. As for KAIST dataset, we use log-average miss rate MR over the false positive per image (FPPI) with the range of [$10^{-2}$, $10^{0}$] to evaluate the pedestrian detection performance.
Settings. All the experiments are conducted with NVIDIA GeForce RTX 3090 GPUs. We implement our framework on the MMDetection library and use the Cascade R-CNN [47] as the basic framework to perform RGB-IR object detection. The detector is trained with an initial learning rate of 2$\times$10-4 for 48 epochs. The batch size is set to 16, and the AdamW [72] optimizer is employed with a weight decay of 0.1. Horizontal flipping is also used for data augmentation.
Results on FLIR and LLVIP datasets. We compare our method with five common mono-modality methods and four competitive multi-modality methods. As shown in Table I, it can be seen that most of the multi-modality detectors are even worse than mono-modality detectors (eg. Faster R-CNN in IR modality). Since the RGB feature interfere with infrared feature under limited illumination conditions, it has a negative impact on the fused features utilized for object detection tasks. However, our UniRGB-IR can effectively solve this problem through the SFI module, enabling the detector to achieve better classification and localization processes.
Results on KAIST dataset. The quantitative results of the different methods on the KAIST dataset are shown in Table II. The experiments are conducted under ‘All-dataset’ settings [53]. We compare our UniRGB-IR with thirteen multi-modal object detection methods. Our model achieves the best performance on the ‘All’, ‘Day’, and ‘Night’ conditions and the other four of five subsets (‘Near’, ‘Far’, ‘None’, ‘Partial’ and ‘Heavy’), and rank second in the ‘Medium’ subset. Furthermore, our detector surpasses the previous best competitor C2Former by 3.18% on the ‘All’ condition, which indicates UniRGB-IR is robust to the complex scenes. Intuitively, we draw the log-average Miss Rate (MR) over the False Positive Per Image (FPPI) curve of these detectors as shown in Figure 5. The above results verify the superiority of our UniRGB-IR.
TABLE IV: Results on the MFNet dataset. The best results are highlighted in red and the second-best are highlighted in blue.
“–” indicates that the author did not the corresponding results.
Methods Unlabeled Car Person Bike Curve Car Stop Guardrail Color Cone Bump mAcc mIoU
Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU
MFNet [18] 98.7 96.9 77.2 65.9 67.0 58.9 53.9 42.9 36.2 29.9 12.5 9.9 0.1 0.0 30.3 25.2 30.0 27.7 45.1 39.7
RTFNet [61] 99.4 98.5 93.0 87.4 79.3 70.3 76.8 62.7 60.7 45.3 38.5 29.8 0.0 0.0 45.5 29.1 74.7 55.7 63.1 53.2
PSTNet [63] 98.8 97.4 89.4 77.0 74.8 67.2 65.8 50.8 42.5 35.2 34.6 25.2 0.0 0.0 32.7 29.9 61.1 55.0 55.5 48.6
FuseSeg [73] 99.0 97.6 93.1 87.9 81.4 71.7 78.5 64.6 68.4 44.8 29.1 22.7 63.7 6.4 55.8 46.9 66.4 47.9 70.6 54.5
ABMDRNet [74] 98.6 97.8 94.3 84.4 90.0 69.6 75.7 60.3 64.0 45.1 44.1 33.1 31.0 5.1 61.7 47.4 66.2 50.0 69.5 54.8
MFFENet [75] 99.3 97.8 91.4 87.1 82.6 74.4 76.7 61.3 58.7 45.6 44.9 30.6 60.0 5.2 64.4 57.0 72.7 40.5 72.3 55.5
EGFNet [67] 98.7 98.0 95.8 87.6 89.0 69.8 80.6 58.8 71.5 42.8 48.7 33.8 33.6 7.0 65.3 8.3 71.1 47.1 72.7 54.8
MTANet [76] 98.4 98.0 95.8 88.1 90.9 71.5 80.3 60.7 75.3 40.9 62.8 38.9 38.7 13.7 63.8 45.9 70.8 47.2 75.2 56.1
MLFNet [77] - 97.3 - 82.3 - 68.1 - 67.3 - 27.3 30.4 - 15.7 - 55.6 - 40.1 - 53.8
FEANet [78] 99.3 98.2 93.3 87.8 82.7 71.1 76.7 61.1 65.5 46.5 26.6 22.1 70.8 6.6 66.6 55.4 77.3 48.9 73.2 55.3
GMNet [79] 99.2 97.5 94.1 86.5 83.0 73.1 76.9 61.7 59.7 44.0 55.0 42.3 71.2 14.5 54.7 48.7 73.1 47.4 74.1 57.3
CCFFNet [30] 98.8 98.3 94.5 89.6 83.6 74.2 73.2 63.1 67.2 50.5 38.7 31.9 30.6 4.8 55.2 49.7 72.9 56.3 68.3 57.6
CCAFFMNet [80] 97.4 95.2 95.0 88.6 86.1 72.9 82.1 67.1 71.2 45.9 32.1 24.8 57.1 17.8 58.0 50.1 76.2 57.8 72.8 57.8
CMX  [31] - 98.3 - 89.4 - 74.8 - 64.7 - 47.3 - 30.1 - 8.1 - 52.4 - 59.4 - 58.2
FDCNet [65] 98.8 98.2 94.1 87.5 91.4 72.4 78.1 61.7 70.1 43.8 34.4 27.2 61.5 7.3 64.0 52.0 74.5 56.6 74.1 56.3
ECGFNet [81] 99.3 97.5 89.4 83.5 85.2 72.1 72.9 61.6 62.8 40.5 44.8 30.8 45.2 11.1 57.2 49.7 65.1 50.9 69.1 55.3
LASNet [82] 97.6 97.4 94.9 84.2 81.7 67.1 82.1 56.9 70.7 41.1 56.8 39.6 59.5 18.9 58.1 48.8 77.2 40.1 75.4 54.9
UniRGB-IR (Ours) 98.3 97.2 94.0 83.7 88.7 64.9 88.0 69.8 53.3 36.8 58.5 41.0 69.7 36.7 77.6 56.2 55.2 47.3 75.7 59.3
TABLE V: Quantitative comparison of RGB-IR salient object detection task on VT821, VT1000 and VT5000 datasets. * represents RGB-D SOD transformed into RGB-T SOD. The best results are highlighted in red and the second-best are highlighted in blue.
Model VT821 VT1000 VT5000
[…]
MMCI* [83] 0.763 0.784 0.618 0.087 0.886 0.892 0.803 0.039 0.827 0.859 0.714 0.055
TANet* [84] 0.818 0.852 0.717 0.052 0.902 0.912 0.838 0.030 0.847 0.883 0.754 0.047
S2MA* [85] 0.811 0.813 0.709 0.098 0.918 0.912 0.848 0.029 0.853 0.864 0.743 0.053
JLDCF* [86] 0.839 0.830 0.726 0.076 0.912 0.899 0.829 0.030 0.861 0.860 0.739 0.050
MTMR [87] 0.725 0.815 0.662 0.109 0.706 0.836 0.715 0.119 0.680 0.795 0.595 0.114
M3S-NIR [88] 0.723 0.859 0.734 0.140 0.726 0.827 0.717 0.145 0.652 0.780 0.575 0.168
SGDL [89] 0.765 0.847 0.731 0.085 0.787 0.856 0.764 0.090 0.750 0.824 0.672 0.089
FMSF [90] 0.760 0.796 0.640 0.080 0.873 0.899 0.823 0.037 0.814 0.864 0.734 0.055
MIDD [91] 0.871 0.895 0.803 0.033 0.915 0.933 0.880 0.027 0.868 0.896 0.799 0.043
ADF [92] 0.810 0.842 0.717 0.077 0.910 0.921 0.847 0.034 0.864 0.891 0.778 0.048
LSNet [93] 0.877 0.911 0.827 0.033 0.924 0.936 0.887 0.022 0.876 0.916 0.827 0.036
UniRGB-IR (Ours) 0.881 0.895 0.806 0.039 0.939 0.943 0.894 0.018 0.906 0.935 0.849 0.027
IV-B RGB-IR Semantic Segmentation
Datasets. Our semantic segmentation experiments are performed on the two public RGB-IR semantic segmentation datasets: PST900 [63] and MFNet [18]. The PST900 dataset is divided into 597 pairs for training and 288 pairs for testing, containing five categories (background, fire extinguisher, backpack, hand drill, and survivor). The MFNet dataset provides 1,569 pairs and contains nine classes (Car, Curve, Person, Guardrail, Car stop, Bike, Bump, Color cone, and Background). The dataset is divided into three parts: training, verification and testing according to the ratio of 2:1:1.
Metrics. Two metrics are utilized to evaluate the performance of semantic segmentation, namely mean accuracy (mAcc) and mean intersection over union (mIoU). Both metrics are calculated by averaging the ratios of the intersection and union of all categories.
Settings. Similarly, as the RGB-IR object detection task, we incorporate our method into the SETR [94] basic framework and implement it on the MMSegmentation library. The fine-tuning process spins a total of 10K iterations with an initial learning rate of 0.01. We employ the SGD optimizer and set the batch size to 16.
Results. The quantitative results of the different RGB-IR segmentation methods on the PST900 and MFNet dataset are shown in Table III and Table IV respectively. The comparison results show that our model significantly outperforms other methods. Specifically, our model obtains the best performance in terms of both the mACC and mIoU. In PST900 dataset, our model performs competitive performance in the Backpack, Hand-Drill and Survivor, outperforming the second-best methods by 2.3%, 0.6% and 0.4% IoU, respectively. In predicting Color Cone class on MFNet dataset, our model achieves the first and second best in terms of both evaluation metrics, with about 11% Acc higher than that of the second-best. These results show that our UniRGB-IR can also be used flexibly for segmentation tasks without any modification, which validates its universality.
Figure 6: Saliency maps of proposed UniRGB-IR and seven representative SOTA RGB-IR SOD methods in difference scenes. From top to bottom are visualization examples from the VT821, VT1000 and VT5000 datasets respectively.
TABLE VI: Ablation studies of key components on the FLIR and PST900 datasets. The best results are highlighted in bold.
FLIR PST900
MFP SFI mAP $\text{mAP}_{50}$ $\text{mAP}_{75}$ mAcc mIoU
39.6 76.3 35.9 90.1 77.5
41.2 77.9 36.5 91.7 78.8
44.1 81.4 40.2 94.3 82.8
IV-C RGB-IR Salient Object Detection
Datasets. Our salient object detection (SOD) experiments are performed on the three public datasets: VT821 [87], VT1000 [89] and VT5000 [92]. The VT821 dataset includes 821 registered RGB and IR images. The VT1000 dataset contains 1000 registered RGB-IR images with simple scenes and aligned images. The VT5000 dataset is a recent large-scale RGB-IR dataset, including a full-day scene under various limited light conditions. As usual in [92], we utilize 2500 image pairs in the VT5000 dataset as the training dataset, and the remaining image pairs along with the image pairs from the VT821 and VT1000 datasets are used as the test datasets.
Metrics. Four metrics are utilized to evaluate the performance of salient object detection namely F-measure (adpF $\uparrow$), E-Measure (adpE $\uparrow$), S-Measure (S $\uparrow$) and Mean absolute error (MAE $\downarrow$). $\uparrow$ and $\downarrow$ denote the higher the better and the lower the better, respectively.
TABLE VII: Ablation of adding SFI module to different stages.
Different Stages with SFI module mAP $\text{mAP}_{50}$ $\text{mAP}_{75}$
Stage 1 Stage 2 Stage 3 Stage 4
41.7 80.0 37.2
43.3 81.6 39.7
44.1 81.4 40.2
42.7 79.9 38.2
TABLE VIII: Ablation of the different attention mechanisms.
Attention Mechanism Complexity mAP $\text{mAP}_{50}$ $\text{mAP}_{75}$
Global Attention [95] Quadratic 36.1 72.8 30.6
Pale Attention [43] Linear 31.6 68.4 24.6
Deformable Attention [44] Linear 44.1 81.4 40.2
Settings. As same as the RGB-IR semantic segmentation task, we incorporate our method into the SETR basic framework and also implement it on the MMSegmentation library. The fine-tuning process spins a total of 10K iterations with an initial learning rate of 0.01. We use the SGD optimizer and set the batch size to 64. For convenience, all input images are resized to 224 $\times$ 224 for testing.
Quantitative Comparison. Table V reports the quantitative comparison results. As can be seen from Table V, our UniRGb-IR outperforms SOTA methods both on VT1000 and VT5000 datasets in all evaluation metrics. Specifically, the $S$, $adpE$, $adpF$ and $MAE$ matrics of our UniRGB-IR achieve 0.906, 0.935, 0.849 and 0.027 on VT5000, all of which are higher than the previous competitor LSNet [93]. These remarkable results clearly indicate that the saliency maps predicted by UniRGB-IR are close to the corresponding ground-truths, which verifies the effectiveness of our method.
Qualitative Evaluation. Qualitative RGB-IR SOD comparisons are shown in Figure 6. The first and third rows show scenes containing small objects. The second and fifth rows show the salient object with the complex backgrounds. All these scenes are suitably handled by our UniRGB-IR, which verifies that our UniRGB-IR is more effective and robustness while providing high-quality results with appropriate detection capabilities.
IV-D Ablation Study
Ablation for components. To investigate the contribution of the SFI and MFP modules, we gradually add each module to the baseline model as shown in Table VI. We stack RGB and IR images and feed them into the Cascade R-CNN detector and SETR framework (ViT backbone) for full fine-tuning on the corresponding tasks, which are considered as our baseline models. Then, we freeze the ViT model and introduce the MFP module to extract contextual multi-scale features (element-wise addition), which results in 1.6% mAP and 1.3% mIoU improvements. Finally, we replace the element-wise addition operation with the SFI module and further improve the mAP and mIoU metrics by 2.9% and 4.0% respectively, which achieve the best performance on both datasets.
Figure 7: Visualization of intermediate results. The $\boldsymbol{F_{mfp}}$ and $\boldsymbol{F_{sfi}}$ features from the first stage are visualized in the third and fourth columns. The tSNE visualizations are also shown in the last two columns.
Figure 8: Training efficiency analysis on the FLIR dataset.
SFI module at different stages. We add the SFI module to the ViT pre-trained model at the beginning of different stages. From Table VII, we can find that by adding the SFI module at the first stage, the detector achieves 41.7% mAP on FLIR dataset. After adding the SFI module in the second and third stages, the performance further improved by about 2% and 3% mAP, respectively. However, continuing to add it to the final stage will reduce the detection performance while also increasing computational overhead. Therefore, we add the SFI module from the first stage to the third stage of ViT model.
Attention type in SFI module. Since the attention mechanism in our SFI module is replaceable, we adopt three popular attention mechanisms in our UniRGB-IR to discuss their impact on model performance. As shown in Table VIII, the detector achieves the best performance with linear complexity by utilizing deformable attention. Thus, the deformable attention is more suitable for our framework and utilized as the default configuration. It is worth noting that it can be replaced by other attention mechanisms to further achieve superior performance.
IV-E Visualization Analysis
[…]
Training efficiency. We further plot the mAP curves for each epoch of UniRGB-IR with different training paradigm to demonstrate the efficiency of UniRGB-IR, as shown in Figure 8. During the training process, all hyperparameters of the two models are the same. From Figure 8, we can find that the convergence speed of the adapter tuning paradigm surpasses that of the full fine-tuning strategy. Moreover, by utilizing the adapter tuning paradigm, our UniRGB-IR achieves superior performance with a smaller number of trainable parameters (about 10% of the full fine-tuning model). The above results verify the efficiency of our method.
V Conclusion
In this paper, we proposed an efficient and scalable framework (named UniRGB-IR) to unify RGB-IR downstream tasks. Our framework contains a Multi-modal Feature Pool module and a Supplementary Feature Injector module. The former extracts contextual multi-scale features from two modality images, and the latter adaptively injects the features into the transformer model. These two modules can be efficiently optimized to complement the pre-trained foundation model with richer RGB-IR features. To evaluate the effectiveness of our method, we incorporated the ViT-Base model into the framework as the pre-trained foundation model and performed various RGB-IR downstream tasks. Extensive experiments verify that our UniRGB-IR can be effectively leveraged as a unified framework for RGB-IR downstream tasks. We believe that our method can be applied to more multi-modal real-world applications.